# Tania Chakraborty, Eylon Caplan, Dan Goldwasser

### v1

@@ 5.3 VSP Inference Analysis
In order to further investigate how these errors led to misclassification, we mapped the clusters onto the full, un-annotated space of visual cues used for the emotion prediction task. To do this, we used the same cluster keywords and the same cosine similarity search mechanism to retrieve the top 500 visual cues for each cluster. We then computed the precision and recall for each emotion, for each cluster. The result for ‘Joy’ can be seen in Figure 6 (other emotions in App. C). We plot $1-P$ and $1-R$ on the y-axis to emphasize bigger errors and visualize their compositions.
We see that for ‘Joy’, recall failures substantially exceed precision failures across every cluster, indicating that the VLM more often misses true visual cues than it hallucinates spurious ones. Particularly, subtle facial movements such as “Brow Furrows” and “Gaze & Eye Behavior” exhibit the highest total failure rates, with nearly all true instances going undetected (very low recall). In sharp contrast, the ‘Smile/Laughter” cluster is almost perfectly handled (both precision and recall failures < 0.1), showing the model’s strength on overt expressions like smiles. Across the board, misinterpretation comprises the largest slice of the error budget, followed by hallucination, while other error types remain minimal. These results suggest that, to improve emotion-prediction accuracy, future work should focus on bolstering the model’s sensitivity to subtle nonverbal signals and reducing its tendency to misinterpret correctly detected cues.
@@ 6 Conclusion
7 Limitations
Variety in Vision models chosen: In this paper we limited our analyses to 2 families of VLMs up to size 26B. More extensive analysis including more models and architectures would lead to better insights.
@@ Appendix C Failure Analysis
Figures 910, and  11 show the error breakdowns for the remaining emotions. Emotions of disgust and surprise had too few data points to contain significant results. Among the figures, clusters with too few points to be significant are also excluded.
@@ Appendix D Prompts
Figure 22: LLM Prompt to classify emotion with Text + Vision.

### v2

@@ 5.3 VSP Inference Analysis
In order to further investigate how these errors led to misclassification, we mapped the clusters onto the full, un-annotated space of visual cues used for the emotion prediction task. To do this, we used the same cluster keywords and the same cosine similarity search mechanism to retrieve the top 500 visual cues for each cluster. We then computed the precision and recall for each emotion, for each cluster. The result for ‘Joy’ can be seen in Figure 6 (other emotions in App. C). We plot 1P1-P and 1R1-R on the y-axis to emphasize bigger errors and visualize their compositions.
We see that for ‘Joy’, recall failures substantially exceed precision failures across every cluster, indicating that the VLM more often misses true visual cues than it hallucinates spurious ones. Particularly, subtle facial movements such as “Brow Furrows” and “Gaze & Eye Behavior” exhibit the highest total failure rates, with nearly all true instances going undetected (very low recall). In sharp contrast, the ‘Smile/Laughter” cluster is almost perfectly handled (both precision and recall failures < 0.1), showing the model’s strength on overt expressions like smiles. Across the board, misinterpretation comprises the largest slice of the error budget, followed by hallucination, while other error types remain minimal. These results suggest that, to improve emotion‐prediction accuracy, future work should focus on bolstering the model’s sensitivity to subtle nonverbal signals and reducing its tendency to misinterpret correctly detected cues.
@@ 6 Conclusion
7 Limitations and Future Work
Variety in Vision models chosen: In this paper we limited our analyses to 3 families of VLMs up to size 26B. The VIBE dataset was tested on 7 models of varying families and sizes. More extensive analysis including more models and architectures could lead to more insights.
@@ 7 Limitations and Future Work
Future Work The limitations mentioned above give rise to some natural directions for future work. It would be worthwhile to study more models on this family of tasks, with an emphasis on architecture and training data to see if there are significant performance differences. Additionally, given the subjective nature of the task, a promising extension of this paper would be to model variations in human judgments as uncertainty estimation of the models. Finally, this paper is a first step into investigating the alignment of multi-modal models and human perceptions of social scenarios; We believe that further investigations and subsequent improvements to these models will lead to socially competent AI systems.
@@ 8 Ethics Statement
Acknowledgments
We thank the reviewers for their insightful comments that helped improve the paper. This work was supported by NSF CAREER award IIS2048001 and the DARPA CCU program. The contents are those of the author(s) and do not necessarily represent the official views of, nor an endorsement by, DARPA, or the US Government.
@@ Appendix A Diagnostic Results
A.1 Experiment to support toggling Levels of VSP Inference
In order to generate output at the correct level of VSP Inference, prompts were manually tuned with care. Additionally we did an experiment to confirm that the outputs generally matched the expected VSP Inference Level; We extracted the top 20 n-grams (n=5, 7, 10) from each of the 3 levels. These were manually inspected (and 10 are provided below in Table5. Additionally, the 20 10-grams were provided to GPt o4-mini-high to compare them with shuffled level order using the prompt:
I’m going to give you top phrases from three different sources. Summarize the main differences between the 3 sources. Output a table where each row is a source. The columns should be “Focus” and “Characteristic Phrases”. Source x: <top n-grams from level 1> Source y: <top n-grams from level 5> Source z: <top n-grams from level 3>
Figure 7: Prompt to confirm VSP Level.
The summary result (re-ordered for readability):
Level 1 (Source: X)
Focus: Detailed, low-level facial movements
Characteristic Phrases:
– “his eyebrows move up and down”
– “her mouth and lips open and close as she speaks”
Level 3 (Source: Z)
Focus: Broader body-language/facial cues
Characteristic Phrases:
– “exhibits a range of body language cues that suggest she is”
– “his eyebrows are slightly furrowed indicating concern”
Level 5 (Source: Y)
Focus: Explicit emotion labeling/classification
Characteristic Phrases:
– “emotions such as joy sadness fear disgust surprise or anger”
– “he is feeling joy. he is smiling and appears to be”
Table 4: The Focus and Characteristic Phrases confirm that the toggling worked as expected.
10-gram Freq 7-gram Freq 5-gram Freq
Level 1 (Detailed, low-level facial movements)
changes in his facial features the video. his eyebrows move 1031 mouth and lips open and close as 2117 eyebrows move up and down 3621
and her mouth and lips open and close as she 1020 maintains a relatively static posture the video. 1975 the the man in the 2690
her mouth and lips open and close as she speaks. 1005 changes in his facial features the video. 1294 mouth and lips open and 2482
changes in her facial features the video. her eyebrows move 998 in his facial features the video. his 1294 and lips open and close 2473
facial features the video. his eyebrows move up and down 975 his facial features the video. his eyebrows 1294 there are no significant changes 2433
in his facial features the video. his eyebrows move up 960 changes in her facial features the video. 1180 are no significant changes in 2433
his facial features the video. his eyebrows move up and 960 in her facial features the video. her 1178 lips open and close as 2128
and his mouth and lips open and close as he 855 her facial features the video. her eyebrows 1178 maintains a relatively static posture 2040
in her facial features the video. her eyebrows move up 854 eyebrows move up and down her forehead 1144 a relatively static posture the 2007
her facial features the video. her eyebrows move up and 854 her mouth and lips open and close 1142 relatively static posture the video. 1975
Level 3 (Broader body-language/facial cues)
a range of body language cues that suggest she is 647 exhibits a range of facial expressions the 2150 a range of facial expressions 3760
exhibits a range of body language cues that suggest she 644 a range of facial expressions the video. 1798 exhibits a range of facial 3704
exhibits a range of facial expressions that suggest he is 601 exhibits a range of facial expressions that 1522 is feeling a mix of 3547
that suggest she is feeling a mix of emotions. her 544 exhibits a range of body language cues 1488 the a man in a 2297
exhibits a range of body language cues that suggest he 513 a range of body language cues that 1482 he is engaged in a 2285
a range of body language cues that suggest he is 513 range of body language cues that suggest 1430 range of facial expressions the 2192
that suggest he is feeling a mix of emotions. his 512 a range of facial expressions that suggest 1332 is engaged in a serious 2036
his eyebrows are slightly furrowed indicating concentration or concern. his 501 his body language suggests that he is 1156 body language cues that suggest 1941
eyebrows are slightly furrowed indicating concentration or concern. his mouth 482 shirt exhibits a range of facial expressions 938 be feeling a mix of 1891
are slightly furrowed indicating concentration or concern. his mouth is 482 the slight furrow in his brow and 936 that he might be feeling 1831
Level 5 (Explicit emotion labeling)
emotions such as joy sadness fear disgust surprise or anger. 680 the in the appears to be feeling 2188 the in the is the 5982
the in the appears to be expressing a mix of 510 the in the appears to be expressing 1446 in the appears to be 4947
strong emotions such as joy sadness fear disgust surprise or 440 joy sadness fear disgust surprise or anger. 1345 the in the appears to 4808
the in the appears to be expressing a sense of 406 appears to be engaged in a conversation 1335 to be engaged in a 3396
such as joy sadness fear disgust surprise or anger. based 404 in the is the man in the 1247 in the is the man 3015
as joy sadness fear disgust surprise or anger. based on 404 the in the is the man in 1244 the appears to be feeling 2554
that he is feeling joy. he is smiling and appears 383 appears to be expressing a mix of 1226 appears to be engaged in 2534
based on these observations it seems that the is feeling 380 to be engaged in a conversation with 1054 in the is the woman 2104
a positive and happy emotion. the in the is the 349 in the appears to be expressing a 963 appears to be expressing a 2090
is feeling joy. he is smiling and appears to be 329 the in the is the man wearing 946 her body language including her 1685
Table 5: Representative n-grams (10, 7, and 5) across Levels 1, 3, and 5.
A.2 Complete VLM-LLM Agent Experiment Results
Here are the f1-scores of the VLM-LLM agent setting for both languages (and overall). Chinese scores are generally slightly higher than the Egnlish ones, which could be due to quirks of the VLM model or the data itself.
Modality Overall Chinese English
T 0.5266 0.5673 0.5006
TV_1 0.5224 0.5582 0.4991
TV_3 0.5199 0.5376 0.5091
TV_5 0.5114 0.5422 0.4952
V_1 0.3002 0.2326 0.3438
V_3 0.4559 0.4568 0.4583
V_5 0.4961 0.5260 0.4805
Table 6: Weighted F1-scores for LLM operating over various modalities. Visual Cues are generated by a VLM given videos of the conversation.
T: Transcript, TV: Transcript & Visual Cues, V: Visual Cues
A.3 Weighted Voting Algorithm Details and Full Results
@@ Appendix C Failure Analysis
Figures 1011, and  12 show the error breakdowns for the remaining emotions. Emotions of disgust and surprise had too few data points to contain significant results. Among the figures, clusters with too few points to be significant are also excluded.
@@ Appendix D Prompts
Figure 23: LLM Prompt to classify emotion with Vision.